Imagine that stock price of a company called ROARJAGUAR (ticker: RJAG), initially a $40 stock, rises to $80.

a) what % is the gain? This gain is called a two dagger (doubling).

b) what percentage is the gain if the stock triples?

c) what percent is the gain if the stock quadruples?
Answer #1

a.

Gain % = ((Final price - Intial Price)/Intial price)*100

b.
Final price = 3*40 = 120

c
Final price = 4*40 = 160